# A Comprehensive Reanalysis of K2-18 b’s JWST NIRISS+NIRSpec Transmission Spectrum

### v1

@@ Abstract
Sub-Neptunes are the most common type of planet in our galaxy. Interior structure models suggest that the coldest sub-Neptunes could host liquid water oceans underneath their hydrogen envelopes—sometimes called “hycean” planets. JWST transmission spectra of the 250 K sub-Neptune K2-18 b were recently used to report detections of CH4 and CO2, alongside weaker evidence of (CH3)2S (dimethyl sulfide, or DMS). Atmospheric CO2 was interpreted as evidence for a liquid water ocean, while DMS was highlighted as a potential biomarker. However, these notable claims were derived using a single data reduction and retrieval modeling framework, which did not allow for standard robustness tests. Here we present a comprehensive reanalysis of K2-18 b’s JWST NIRISS SOSS and NIRSpec G395H transmission spectra, including the first analysis of the second-order NIRISS SOSS data. We incorporate multiple well-tested data reduction pipelines and retrieval codes, spanning 60 different data treatments and over 250 atmospheric retrievals. We confirm the detection of CH4 (), with a volume mixing ratio of , but we find no statistically significant or reliable evidence for CO2 or DMS. Finally, we quantify the observed atmospheric composition using photochemical-climate and interior models, demonstrating that our revised composition of K2-18 b can be explained by an oxygen-poor mini-Neptune without requiring a liquid water surface or life.
@@
exoTEDRF (Feinstein et al., 2023; Radica et al., 2023),
@@ 1 Introduction
While JWST transmission spectra have succeeded in probing the giant exoplanet atmospheres, decisive results have been more elusive for the small and cold ends of the exoplanet population. For super-Earths and terrestrial worlds (), JWST transmission spectra largely rule out thick H2-dominated atmospheres (e.g., Lustig-Yaeger et al. 2023; Lim et al. 2023; Alderson et al. 2024; Radica et al. 2024a) or are degenerate with unocculted stellar active regions (e.g., Moran & Stevenson et al. 2023; May & MacDonald et al. 2023). However, sub-Neptunes111We use “sub-Neptune” when categorizing planets by size (c.f. super-Earth), and “mini-Neptune” when categorizing planets by interior structure (c.f. hycean)., planets with radii , likely possess significant atmospheres dominated by relatively light molecules (e.g., H2 or H2O), which aids their atmospheric detectability. While searches for atmospheres on several sub-Neptunes have yielded non-detections (e.g. Wallack et al., 2024) or weak evidence (e.g., Damiano et al., 2024; Cadieux et al., 2024), an emerging population now has definitive atmospheric detections. To date, JWST has detected atmospheres on four sub-Neptunes: GJ 3470 b (Beatty et al., 2024), GJ 9827 d (Piaulet-Ghorayeb et al., 2024), TOI-270 d (Benneke et al., 2024; Holmberg & Madhusudhan, 2024), and K2-18 b (Madhusudhan et al., 2023). Of these planets, the cold ( 250 K) sub-Neptune K2-18 b has attracted significant observational and theoretical interest in its atmospheric composition and interior structure.
@@
The K2-18 system was identified as a potential transiting exoplanet system in the first campaign of NASA’s K2 mission (Montet et al., 2015). Further transit observations by Spitzer (Benneke et al., 2017) and radial velocity follow-up (Cloutier et al., 2017; Sarkis et al., 2018) confirmed K2-18 b as a habitable zone planet, with a refined radial velocity extraction method further improving its mass precision (Radica et al., 2022a). The system is now known to host at least two planets (Cloutier et al., 2019; Radica et al., 2022a), though only K2-18 b transits its early M dwarf host star. The combination of its cool temperature (255 K, assuming an albedo of 0.3; Benneke et al. 2019), relatively large planetary radius (2.61 ), and small stellar radius (0.4445 ) has led to K2-18 b being one of the best targets for the atmospheric characterization of a potentially habitable exoplanet.
@@
Initial transmission spectroscopy observations of K2-18 b with the Hubble Space Telescope (HST) yielded a detection of a H2-dominated atmosphere, with evidence of either gas phase H2O or CH4. The HST Wide Field Camera 3 (WFC3) transmission spectrum of K2-18 b was initially interpreted as showing a detection of H2O (Tsiaras et al., 2019; Benneke et al., 2019; Madhusudhan et al., 2020). However, later work by Barclay et al. (2021) argued that the low-resolution WFC3 data could alternatively be explained by CH4, due to the similar shapes of H2O and CH4 absorption over the 1.1–1.7 WFC3 bandpass. These competing interpretations of K2-18 b’s atmospheric composition resulted in an inability to definitively constrain the nature of its atmosphere and interior through HST observations alone.
@@
The astrobiological assessment of H2-dominated sub-Neptunes such as K2-18 b is at an early stage, and hence any claims of life must be tempered with extreme caution. Madhusudhan et al. (2023) highlighted a low-significance inference of DMS ( from their best-fitting model) as potential evidence for life, given DMS is significantly produced by marine phytoplankton on Earth (Charlson et al., 1987; Barnes et al., 2006). However, the recent detection of DMS on a comet by Hänni et al. (2024) demonstrates that DMS can also be produced in abiotic environments. Further, Tsai et al. (2024) demonstrated that a DMS flux equivalent to 20 times that produced by Earth’s marine life would be required for DMS to reach detectable levels on a hycean planet with K2-18 b-like conditions. Recently, Glein (2024) recently showed that CO2-consuming and CH4-producing life is energetically viable on K2-18 b, if one assumes a hycean scenario and the atmospheric abundances reported in Madhusudhan et al. (2023).
@@
The plausibility of the hycean interpretation for K2-18 b relies critically on the retrieved chemical abundances reported by Madhusudhan et al. (2023). However, to date, there has been no independent analysis of the original JWST transmission spectrum of K2-18 b. The gold standard for interpreting JWST spectra of exoplanet atmospheres is to apply multiple data reduction techniques and retrieval codes to confirm the reproducibility and robustness of the results — this philosophy has been broadly adopted by the exoplanet community (e.g., Coulombe et al., 2023; Taylor et al., 2023; Grant et al., 2023; Powell et al., 2024; Gressier et al., 2024; Banerjee et al., 2024; Piaulet-Ghorayeb et al., 2024; Welbanks et al., 2024). However, the K2-18 b analysis by Madhusudhan et al. (2023) investigated only a single reduction and retrieval model framework. Performing several reductions on the same data set quantifies the impact of choices such as outlier rejection, noise correction, and limb darkening treatments that occur when using independent pipelines. Similarly, performing retrievals with different codes quantifies how model-level assumptions (e.g., opacity sources, aerosol parameterization) propagate into atmospheric inferences. Atmospheric inferences that withstand these tests are considered robust. On the other hand, severe inconsistencies (e.g., a molecular detection seen only with one reduction), demonstrate that the statistical evidence for such an atmospheric inference must be carefully assessed. Here, we provide the first comprehensive analysis of K2-18 b’s JWST NIRISS SOSS and NIRSpec G395H transmission spectra, using multiple data reduction and retrieval codes, to offer an updated assessment of K2-18 b’s atmospheric composition and possible interior structures.
@@ 2 NIRISS Data Reductions
The K2-18 system was observed by JWST’s Near Infrared Imager and Slitless Spectrograph (NIRISS; Doyon et al., 2012, 2023) instrument using the Single Object Slitless Spectroscopy (SOSS; Albert et al., 2023) mode from 13:49:20 UTC to 19:36:05 UTC (for a total of 4.9 hours) on June 1, 2023 as part of JWST GO Program 2722 (PI: N. Madhusudhan). The observation, which covered one transit of K2-18 b, used the GR700XD grism, CLEAR filter, SUBSTRIP256 subarray, and NISRAPID readout pattern. The exposure consisted of 648 integrations with 4 groups per integration, and there were no mirror tilt events or high-gain antenna movements during the observation. The first spectral order yielded a wavelength coverage of 0.85–2.85 m which, when combined with the second order’s wavelength coverage of 0.6–1.4 m, results in a total uninterrupted wavelength coverage of 0.6–2.85 m.
We perform two reductions with two independent pipelines: FIREFLy and exoTEDRF222We do not perform NIRISS reductions with the Eureka! pipeline, as support for this instrument mode is currently pending.. To allow a transparent comparison between our reduction approaches and Madhusudhan et al. (2023) (who used the JExoRES pipeline Holmberg & Madhusudhan 2023), we provide an overview of the most important reduction-level configuration choices for NIRISS in Table 1 and show additional diagnostic plots for our reductions in Appendix C.
@@
| Wavelength calibration | pastasoss | Cross correlation with PHOENIX | ? |
@@
| Error bar inflation | Red noise via Allan variance | Additive error inflation | ? |
@@
Note. — “Default” refers to the default handling of the step in the jwst pipeline. The system parameters (, , and ) are consistent across each reduction, though they were calculated in Madhusudhan et al. (2023) based on their reduction up to that point. This is not an exhaustive list of reduction steps. ‘?’ indicates reduction configuration settings not mentioned in Madhusudhan et al. (2023).
@@ 2.1 FIREFLy
We use the FIREFLy pipeline (Rustamkulov et al., 2022, 2023), which has been recently updated to better support NIRISS/SOSS observations (e.g., Liu & Wang et al. in prep, Wang et al. in prep), to re-reduce the K2-18 NIRISS/SOSS observations. We start from the raw, uncalibrated files on MAST and followed the jwst pipeline for group-level and instrument corrections. We start by performing the saturation and superbias steps, but we skip the dark current step. To correct the 1/ noise at the group level, we first subtract the zodiacal background at each group. This is done by scaling the flux jump caused by the reflection of zodiacal light off the pick-off mirror in the STScI SOSS background model to the flux jump present in the observed data and adding a constant offset to match the flux level of the model SOSS background before and after the jump to the observed data. We then mask out the bright spectral trace and notable zeroth order contaminants present in the frame and performed 1/ destriping by subtracting the median of each PSF-masked and background-subtracted column. In addition, we subtract each image from the temporal median, which we create as a running median of a 7 frame window, to reveal any remaining 1/ noise and subtract it. This removes high frequency structure while preserving the slowly changing source flux. Because the background levels are real counts and not the detector’s bias level, we add the background back after subtracting the 1/ noise in order to ensure that the jwst pipeline’s bias correction and up-the-ramp fitting procedures are unaffected. We perform the default linearity and ramp fitting step, but we skip the jump step. For the wavelength calibration and order tracing, we use pastasoss (Baines et al., 2023), which predicts the GR700XD PSF positions and the wavelength solutions accordingly with sub-pixel accuracy based on the position of pupil wheels. We apply an integration-level 1/ subtraction and clean outlier bad/hot pixels.
@@
We then perform Stage 3, FIREFLy’s stellar and spectral extraction step. We apply a temporal outlier cut, remove the background for the final time, and perform an additional 1/ noise removal at the integration level. Both the background removal and 1/ noise subtraction follow the same algorithm as the group level. To extract the white light curves, we sum all the flux in order 1 and the flux from wavelengths within m for order 2. We align frames and conduct box extractions on each order with aperture widths that minimize the scatter of the white light curve for each order–30 pixels for order 1 and 23 pixels for order 2.
We model the transit light curves with the batman Python package (Kreidberg, 2015). We note a spot-crossing event in the NIRISS light curve, so we trim out those integrations accordingly. Informed by the Bayesian Information Criterion (BIC), we choose the systematics vector to be composed solely of a linear trend for both order 1 and order 2. We fit for the orbital parameters: , the stellar radius-scaled semimajor axis; , the impact parameter; and , the center-of-transit time. We find that our fitted values are consistent with those reported by Madhusudhan et al. (2023); we therefore fix the orbital parameters to theirs for a consistent comparison. We jointly fit the transit light curves and the systematics using the Markov Chain Monte Carlo (MCMC) sampler emcee (Foreman-Mackey et al., 2013), as implemented in the lmfit Python package (Newville et al., 2024). We show the white light curve fits for the NIRISS SOSS orders 1 and 2 separately in Figure 1, with the spot-crossing event highlighted.
For limb darkening, we perform a two-step process. We first fit the white light curve and spectrophotometric limb darkening coefficients, before fixing them to a 3D Stagger grid stellar atmosphere model (Magic et al., 2015) with an offset. We experiment with both the Kipping (2013) {, } and the conventional quadratic and limb darkening parametrizations. Since both approaches result in virtually identical transmission spectra, we adopt {, } for each of our fits (as tends to be close to zero and less informative about the effects of limb darkening on the transit’s shape). Combining the fitted and from order 1 and order 2, we find that these fitted limb-darkening coefficients trace well the limb darkening coefficients from the Stagger grid of 3D stellar atmosphere models, with an offset of for and for . We therefore fix these offset model limb darkening coefficients for our spectropohotometric fit.
We fit the spectroscopic light curves at a two-pixel binning level (the same binning strategy as Madhusudhan et al., 2023) as well as at resolving powers and . We use two-pixel binning for NIRISS SOSS rather than one-pixel binning as its PSF is two pixels wide. As with the white light curve fit, we fix the orbital parameters to those used by Madhusudhan et al. (2023). For the spectroscopic light curve fits, we use the lmfit Python package to perform the Levenberg-Marquardt algorithm to fit the radius ratio for the spectrophotometric light curves as well as a linear systematic term for each bin. We account for red noise by inflating error bars in quadrature for wavelength bins where the Allan variance is above log linear. This does not impact the error bars for most points, but for those that are, it is typically by a factor of no more than two.
@@
For the light curve analysis, we first construct two separate white light curves by summing all the flux in order 1, and wavelengths for order 2. We then jointly fit the white light curves using the flexible exoUPRF library (Radica, 2024) following the procedure in Radica et al. (2024b, a). The fitted light curve model is composed of an astrophysical component (i.e., a batman transit model), and a systematics component. For the astrophysical component, we share the orbital parameters (i.e., time of mid-transit, orbital inclination and scaled semi-major axis, as well as the orbital eccentricity and argument of periastron) between the two orders, but separately fit the scaled planet radius and two parameters of the quadratic limb darkening law to each. The systematics component consists of a linear trend with time, fit independently to each order, as well as a scalar jitter term added in quadrature to the flux errors. Finally, we note a star spot crossing event approximately 45 minutes before mid transit. Instead of commonly-used prescriptions like spotrod (Béky et al., 2014; Madhusudhan et al., 2023; Fournier-Tondreau et al., 2024a) we simply model the spot crossing using a Gaussian profile, which has been shown to produce equally accurate fits to spot crossing events in transit light curves (Roy et al. in prep). We fit the amplitude of the Gaussian independently to each order, but share the width and position of the Gaussian between the two. In all, our white light fit has 19 free parameters, and we use wide, uninformative priors for each parameter. For the fits, we use the affine-invariant sampler emcee with 40 chains and using 50000 steps per chain. We discard the first 80% as burn-in.
We then fit the spectrophotometric light curves at three different resolutions — in bins with widths of two pixels (since the size of the SOSS point spread function is roughly two pixels in the dispersion direction), and constant resolutions of and . For the spectrophotometric fits, we fix the orbital parameters to those used by Madhusudhan et al. (2023) to allow a consistent comparison between the two works. We also test fixing the orbital parameters to the updated solution presented in Radica et al. (2022a), but find there to be negligible changes in the resulting transmission spectra. We also fix the location and width of the Gaussian spot model to our best-fitting white light curve values, but fit the amplitude for each wavelength bin. Finally, we freely fit the quadratic limb darkening parameters, systematic slope, and additive jitter term to each bin, resulting in a total of eight free parameters.

### v2

@@ Abstract
Sub-Neptunes are the most common type of planet in our galaxy. Interior structure models suggest that the coldest sub-Neptunes could host liquid water oceans underneath their hydrogen envelopes—sometimes called “hycean” planets. JWST transmission spectra of the 250 K sub-Neptune K2-18 b were recently used to report detections of CH4 and CO2, alongside weaker evidence of (CH3)2S (dimethyl sulfide, or DMS). Atmospheric CO2 was interpreted as evidence for a liquid water ocean, while DMS was highlighted as a potential biomarker. However, these notable claims were derived using a single data reduction and retrieval modeling framework, which did not allow for standard robustness tests. Here we present a comprehensive reanalysis of K2-18 b’s JWST NIRISS SOSS and NIRSpec G395H transmission spectra, including the first analysis of the second-order NIRISS SOSS data. We incorporate multiple well-tested data reduction pipelines and retrieval codes, spanning 60 different data treatments and over 250 atmospheric retrievals. We confirm the detection of CH4 (), with a volume mixing ratio range , but we find no statistically significant or reliable evidence for CO2 or DMS. Finally, we assess the retrieved atmospheric composition using photochemical-climate and interior models, demonstrating that our revised composition of K2-18 b can be explained by an oxygen-poor mini-Neptune without requiring a liquid water surface or life.
@@
exoTEDRF (Feinstein et al., 2023; Radica et al., 2023; Radica, 2024),
@@ 1 Introduction
While JWST transmission spectra have succeeded in probing the giant exoplanet atmospheres, decisive results have been more elusive for the small and cold ends of the exoplanet population. For super-Earths and terrestrial worlds (), JWST transmission spectra largely rule out thick H2-dominated atmospheres (e.g., Lustig-Yaeger et al. 2023; Lim et al. 2023; Alderson et al. 2024a; Radica et al. 2024a) or are degenerate with unocculted stellar active regions (e.g., Moran & Stevenson et al. 2023 2023; May & MacDonald et al. 2023 2023). However, sub-Neptunes111We use “sub-Neptune” when categorizing planets by size (c.f. super-Earth), and “mini-Neptune” when categorizing planets by interior structure (c.f. hycean)., planets with radii , likely possess significant atmospheres dominated by relatively light molecules (e.g., H2 or H2O), which aids their atmospheric detectability. While searches for atmospheres on several sub-Neptunes have yielded non-detections (e.g. Wallack et al., 2024) or weak evidence (e.g., Damiano et al., 2024; Cadieux et al., 2024), an emerging population now has definitive atmospheric detections. To date, JWST has detected atmospheres on five sub-Neptunes: GJ 3470 b (Beatty et al., 2024), GJ 9827 d (Piaulet-Ghorayeb et al., 2024), TOI-270 d (Benneke et al., 2024; Holmberg & Madhusudhan, 2024), TOI-421 b (Davenport et al., 2025), and K2-18 b (Madhusudhan et al., 2023). Of these planets, the cold ( 250 K) sub-Neptune K2-18 b has attracted significant observational and theoretical interest in its atmospheric composition and interior structure.
@@
The K2-18 system was identified as a potential transiting exoplanet system in the first campaign of NASA’s K2 mission (Montet et al., 2015). Further transit observations by Spitzer (Benneke et al., 2017) and radial velocity follow-up (Cloutier et al., 2017; Sarkis et al., 2018) confirmed K2-18 b as a habitable zone planet, with a refined radial velocity extraction method further improving its mass precision (Radica et al., 2022a). The system is now known to host at least two planets (Cloutier et al., 2019; Radica et al., 2022a), though only K2-18 b transits its early M dwarf host star. The combination of its cool temperature (255 K, assuming an albedo of 0.3; Benneke et al. 2019a), relatively large planetary radius (2.61 ), and small stellar radius (0.4445 ) has led to K2-18 b being one of the best targets for the atmospheric characterization of a potentially habitable exoplanet.
@@
Initial transmission spectroscopy observations of K2-18 b with the Hubble Space Telescope (HST) yielded a detection of a H2-dominated atmosphere, with evidence of either gas phase H2O or CH4. The HST Wide Field Camera 3 (WFC3) transmission spectrum of K2-18 b was initially interpreted as showing a detection of H2O (Tsiaras et al., 2019; Benneke et al., 2019a; Madhusudhan et al., 2020). However, later work by Barclay et al. (2021) argued that the low-resolution WFC3 data could alternatively be explained by CH4, due to the similar shapes of H2O and CH4 absorption over the 1.1–1.7 WFC3 bandpass. These competing interpretations of K2-18 b’s atmospheric composition resulted in an inability to definitively constrain the nature of its atmosphere and interior through HST observations alone.
@@
The astrobiological assessment of H2-dominated sub-Neptunes such as K2-18 b is at an early stage, and hence any claims of life must be tempered with extreme caution. Madhusudhan et al. (2023) highlighted a low-significance inference of DMS ( from their best-fitting model) as potential evidence for life, given DMS is significantly produced by marine phytoplankton on Earth (Charlson et al., 1987; Barnes et al., 2006). Such a DMS flux would have to be equivalent to 20 times that produced by Earth’s marine life to reach detectable levels on a hycean planet with K2-18 b-like conditions (Tsai et al., 2024). Glein (2024) showed that metabolic reactions consuming CO2 and/or DMS and producing CH4 can potentially produce sufficient free energy to sustain a biosphere (if one assumes a hycean scenario and the atmospheric abundances reported in Madhusudhan et al., 2023) on K2-18 b. However, DMS in this metabolic pathway cannot strictly be considered a biosignature due to it being an energy source for methane-producing life rather than a by-product of life. Furthermore, the recent detection of DMS on a comet by Hänni et al. (2024) demonstrates that DMS can also be produced in abiotic environments. These issues with DMS as potential evidence for a biosphere on K2-18 b in combination with the lack of detailed cross section data for it complicate any potential investigation of the planet’s habitability. Indeed, the state-of-the-art line data for DMS (HITRAN; Sharpe et al., 2004) lacks pressure and temperature dependence, potentially biasing retrieval results that include it (Heng & Kitzmann, 2017; Welbanks & Madhusudhan, 2019; Niraula et al., 2022).
@@
The plausibility of the hycean interpretation for K2-18 b relies critically on the retrieved chemical abundances reported by Madhusudhan et al. (2023). However, to date, there has been no independent analysis of the original JWST transmission spectrum of K2-18 b. The gold standard for interpreting JWST spectra of exoplanet atmospheres is to apply multiple data reduction techniques and retrieval codes to confirm the reproducibility and robustness of the results — this philosophy has been broadly adopted by the exoplanet community (e.g., Coulombe et al., 2023; Taylor et al., 2023; Grant et al., 2023; Powell et al., 2024; Gressier et al., 2024; Banerjee et al., 2024; Piaulet-Ghorayeb et al., 2024; Welbanks et al., 2024). However, the K2-18 b analysis by Madhusudhan et al. (2023) investigated only a single reduction and retrieval model framework. Performing several reductions on the same data set quantifies the impact of choices such as outlier rejection, noise correction, and limb darkening treatments that occur when using independent pipelines. Similarly, performing retrievals with different codes can illustrate how some model-level assumptions (e.g., opacity sources, aerosol parameterization) propagate into atmospheric inferences (e.g., Mai & Line, 2019; Welbanks & Madhusudhan, 2019; Barstow et al., 2020; Niraula et al., 2023; Nixon et al., 2024a). Atmospheric inferences that withstand these tests are considered robust. On the other hand, severe inconsistencies (e.g., a molecular detection seen only with one reduction), demonstrate that the statistical evidence for such an atmospheric inference must be carefully assessed. Here, we provide the first comprehensive analysis of K2-18 b’s JWST NIRISS SOSS and NIRSpec G395H transmission spectra, using multiple data reduction and retrieval codes, to offer an updated assessment of K2-18 b’s atmospheric composition and possible interior structures.
We note that during the revision of this study, Madhusudhan et al. (2025) presented the MIRI LRS transmission spectrum of K2-18 b. While Madhusudhan et al. (2025) reported independent evidence for DMS and/or DMDS (dimethyl disulfide) at , subsequent studies have demonstrated that the MIRI data is well-fit by a flat line (Taylor, 2025), that DMS and DMDS do not provide a uniquely good fit relative to other molecules (Welbanks et al., 2025; Pica-Ciamarra et al., 2025), and that DMS or DMDS are not detected when the MIRI data is analyzed in concert with the near-infrared data (Luque et al., 2025). These studies highlight the low information content of the MIRI data relative to the shorter wavelength near-infrared data, which we focus on in this study.
@@ 2 NIRISS Data Reductions
The K2-18 system was observed by JWST’s Near Infrared Imager and Slitless Spectrograph (NIRISS; Doyon et al., 2012, 2023) instrument using the Single Object Slitless Spectroscopy (SOSS; Albert et al., 2023) mode as part of JWST GO Program 2722 (PI: N. Madhusudhan - Madhusudhan, 2023). The details of this observation are described thoroughly in Madhusudhan et al. (2023). We perform two reductions of this data set with two independent pipelines: FIREFLy and exoTEDRF222We do not perform NIRISS reductions with the Eureka! pipeline, as support for this instrument mode is currently pending.. To allow a transparent comparison between our reduction approaches and Madhusudhan et al. (2023) (who used the JExoRES pipeline Holmberg & Madhusudhan 2023), we provide an overview of the most important reduction-level configuration choices for NIRISS in Table 1. Both of our reductions fit the spectroscopic light curves at a two-pixel binning level (the same binning strategy as Madhusudhan et al., 2023) as well as at resolving powers and . We use two-pixel binning for NIRISS SOSS rather than one-pixel binning as its PSF is two pixels wide. Additionally, we show the point-by point discrepancies between reductions in Appendix C, as has been done previously for analyses involving multiple reductions where small point-by-point differences could yield different atmospheric interpretations (e.g., Alderson et al., 2024b; Alam et al., 2025; Louie et al., 2025; Alderson et al., 2025; Kirk et al., 2025).
@@
| Wavelength calibration | pastasoss(Baines et al., 2023) | Cross correlation with PHOENIX | ? |
@@
| Error bar inflation | Red noise via residual RMS-bin | Additive error inflation | ? |
| size trend | |||
@@
Note. — “Default” refers to the default handling of the step in the jwst pipeline. The system parameters (, , and ) are consistent across each reduction, though they were calculated in Madhusudhan et al. (2023) based on their reduction up to that point. The listed orbital parameters were used to either confirm consistency with Madhusudhan et al. (2023) (FIREFLy) or to show negligible differences between resulting spectra with an alternative orbital solution (exoTEDRF, from Radica et al., 2022a). This is not an exhaustive list of reduction steps. ‘?’ indicates reduction configuration settings not mentioned in Madhusudhan et al. (2023).
@@ 2.1 FIREFLy
We use the FIREFLy pipeline (Rustamkulov et al., 2022, 2023), which has been recently updated to better support NIRISS/SOSS observations (Liu & Wang et al. 2025, L. C. Wang et al. in prep), to re-reduce the K2-18 NIRISS/SOSS observations. A more thorough inventory of the steps we use are included in Table 1; here we describe the most pertinent choices made. To correct the 1/ noise at the group level, we first subtract the zodiacal background at each group. This is done by scaling the flux jump caused by the reflection of zodiacal light off the pick-off mirror in the STScI SOSS background model to the flux jump present in the observed data and adding a constant offset to match the flux level of the model SOSS background before and after the jump to the observed data. We then mask out the bright spectral trace and notable zeroth order contaminants present in the frame and performed 1/ destriping by subtracting the median of each PSF-masked and background-subtracted column. In addition, we subtract each image from the temporal median, which we create as a running median of a 7 frame window, to reveal any remaining 1/ noise and subtract it. This removes high frequency structure while preserving the slowly changing source flux. Because the background levels are real counts and not the detector’s bias level, we add the background back after subtracting the 1/ noise in order to ensure that the jwst pipeline’s bias correction and up-the-ramp fitting procedures are unaffected. We also apply an integration-level 1/ subtraction and clean outlier bad/hot pixels.
@@
In Stage 3, we apply a temporal outlier cut, remove the background for the final time, and perform an additional 1/ noise removal at the integration level. Both the background removal and 1/ noise subtraction follow the same algorithm as the group level. To extract the white light curves, we sum all the flux in order 1 and the flux from wavelengths within m for order 2. We align frames and conduct box extractions on each order with aperture widths that minimize the scatter of the white light curve for each order.
Moving on to light curve fitting, we trim out integrations corresponding to a spot-crossing event in the NIRISS light curve. Informed by the Bayesian Information Criterion (BIC), we choose the systematics vector to be composed solely of a linear trend for both order 1 and order 2. We find that our fitted values are consistent with those reported by Madhusudhan et al. (2023); we therefore fix the orbital parameters to theirs for a consistent comparison in both the white light curve and spectrophotometric fits. For limb darkening, we perform a two-step process of first fitting for spectroscopic limb darkening coefficients and then fixing them to an offset stellar atmosphere model. We experiment with both the Kipping (2013) {, } and the quadratic and parametrizations, choosing {, } as in particular tends to be uninformative about light curve shape despite both spectra being virtually indistinguishable. Combining the fitted and from order 1 and order 2, we find that these fitted limb-darkening coefficients trace well the limb darkening coefficients from the Stagger grid of 3D stellar atmosphere models, with an offset of for and for . We show the white light curve fits for the NIRISS SOSS orders 1 and 2 separately in Figure 1, with the spot-crossing event highlighted. We account for red noise in our spectrophotometric fit by inflating error bars in quadrature for wavelength bins where the residual RMS-bin size trend is above log linear (Pont et al., 2006; Winn et al., 2008). This does not impact the error bars for most points, but for those that are, it is typically by a factor of no more than two.
@@
For the light curve analysis, we first construct two separate white light curves by summing all the flux in order 1, and wavelengths for order 2. We then jointly fit the white light curves using the flexible exoUPRF library (Radica, 2024) following the procedure in Radica et al. (2024b, a). The fitted light curve model is composed of an astrophysical component (i.e., a batman transit model), and a systematics component. For the astrophysical component, we share the orbital parameters (i.e., time of mid-transit, orbital inclination and scaled semi-major axis, as well as the orbital eccentricity and argument of periastron) between the two orders, but separately fit the scaled planet radius and two parameters of the quadratic limb darkening law to each. We put Gaussian priors on the orbital parameters based on the findings from Madhusudhan et al. (2023). The systematics component of our light curve fit consists of a linear trend with time, fit independently to each order, as well as a scalar jitter term added in quadrature to the flux errors. Instead of commonly-used prescriptions like spotrod (Béky et al., 2014; Madhusudhan et al., 2023; Fournier-Tondreau et al., 2024a) to model the spot-crossing event, we simply model it with a Gaussian profile, which has been shown to produce equally accurate fits to spot crossing events in transit light curves (Roy et al. in prep). We fit the amplitude of the Gaussian independently to each order, but share the width and position of the Gaussian between the two. In all, our white light fit has 19 free parameters, and we use wide, uninformative priors for each parameter. For the fits, we use the affine-invariant sampler emcee with 40 chains and using 50000 steps per chain. We discard the first 80% as burn-in.
For the spectrophotometric fits, we fix the orbital parameters to the best-fitting values from the white light curve fits. We test fixing the orbital parameters to the updated solution presented in Radica et al. (2022a), but find negligible changes in the resulting transmission spectra. We also fix the location and width of the Gaussian spot model to our best-fitting white light curve values, but fit the amplitude for each wavelength bin. Finally, we freely fit the quadratic limb darkening parameters, systematic slope, and additive jitter term to each bin, resulting in a total of eight free parameters.